# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v1

@@ 4.3 Identification of neural dynamics in mice visual cortex
Since NN-ResDMD directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics. To evaluate how effectively NN-ResDMD reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel ResDMD. These methods are selected as representative approaches for handling high-dimensional data.
@@
The dataset is part of the open dataset on mice from the competition "Sensorium 2023" [26, 27]. In the experiments, mice viewed natural videos while their neural signals were recorded via calcium imaging in the primary visual cortex, reflecting the activity of thousands of neurons. Here, we focus on the state partitioning of neural signals. Specifically, in each mouse, six video stimuli were repeatedly shown, creating ideal conditions to define brain states. Neural activity during repeated trials with the same stimuli is assumed to reflect the same underlying dynamic system, enabling Koopman decomposition methods to uncover and separate these brain states.
@@
We applied NN-ResDMD and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel ResDMD) to these datasets, using different implementations and Koopman subspace dimensions. For NN-ResDMD, we trained dictionaries on all snapshots from each mouse to avoid overfitting, reduced the data to 300 dimensions via SVD, and selected 501 eigenfunctions. The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations. For Hankel-DMD, we built a Hankel matrix with a delay of 50, producing 50 eigenfunctions per trial. In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions. For Kernel ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on [28]. See Appendix A.7.4 for method details and Appendix A.8 for dictionary size justification. These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 9A, and Appendix Figure 10A, are compared to the ground truth trial identities.
@@
The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli. To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, distinguishing them from other states. If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others. This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli. Note that averaged trial differences are even visibly clear for the NN-ResDMD case.
Figure 6: NN-ResDMD outperforms Hankel-DMD in identifying latent dynamic structures in neural signals with a dictionary size of 501. (A) (Top) 500 Koopman eigenfunctions estimated by NN-ResDMD across 6 states characterized by different video stimuli in an example mouse. Each trial contains 300 data points (10s at 50Hz). (Bottom) 50 Koopman eigenfunctions approximated by Hankel-DMD, each 50 points long, reflecting the dimension of the Hankel matrix. (B) 2D representation of Koopman eigenfunctions for all tested mice, computed by NN-ResDMD and reduced via Multidimensional Scaling (MDS). Trials of the same state cluster well. (C) Same as (B) but computed with Hankel-DMD, showing no clear state separation. (D) 2D representation of Koopman eigenfunctions for the first mouse, computed by EDMD with an RBF basis. See Appendix Figure 9 for full results. (E) Same as (D) but computed with Kernel ResDMD. See Appendix Figure 10 for full results. (F) Davies-Bouldin Indices (DBIs) evaluating clustering quality across four methods (NN-ResDMD, Hankel-DMD, EDMD+RBF, and Kernel ResDMD) for five mice. Lower DBI values for NN-ResDMD indicate better clustering.
@@
We use Multi-dimensional Scaling (MDS) to visualize how these eigenfunction-based features cluster according to ground truth states. MDS reduces data dimensionality based on similarities, making it ideal for visualizing clustering performance. While UMAP and t-SNE are alternative methods, we show MDS results in 2D space (Figure 6B-E), with similar results for UMAP and t-SNE in the supplementary materials (Appendix Figure 8, Appendix Figure 9C,D and Appendix Figure 10C,D).
@@
The 2D MDS visualization reveals clear separation of features for all 5 mice using NN-ResDMD (Figure 6B), whereas no other method shows clear clustering (Figure 6C-E, Appendix Figure 9B, Appendix Figure 10B). To quantify this clustering, we calculate the Davies-Bouldin index (DBI), a measure of clustering quality that assesses how compact and well-separated the clusters are. A lower DBI indicates more compact clusters that are farther apart from each other, which corresponds to better clustering. The DBI is significantly lower for NN-ResDMD (Figure 6F), suggesting that it captures the latent dynamic structure more effectively than all three other methods. Similar clustering patterns are confirmed with UMAP and t-SNE (Appendix Figure 11).
@@ 5 Conclusion and future work
Koopman spectral components (eigenpairs) are fundamental to understanding dynamical systems, as they reveal intrinsic patterns and structures underlying complex temporal behavior through a linear framework for analyzing nonlinear dynamics. In this paper, we introduced NN-ResDMD, a method for estimating eigenpairs by minimizing spectral residuals, eliminating ResDMD’s need to filter precomputed results. Despite higher computational costs, using neural networks to learn eigenpairs provides a significant advantage by capturing patterns automatically and reducing manual intervention in basis selection. This flexibility is particularly beneficial for high-dimensional systems where traditional methods often struggle. Our experiments demonstrate that NN-ResDMD outperforms classical methods—including EDMD, Hankel-DMD, ResDMD, and their variants—in uncovering critical spatiotemporal characteristics of nonlinear dynamics.
Despite the advantages, NN-ResDMD has several limitations and we discuss the major ones here. First, the neural network structure incurs higher computational costs compared to classical approaches, making it unsuitable for real-time learning tasks (see a brief discussion in Appendix A.5). Second, the deterministic nature of the framework does not account for stochastic aspects of the system, such as those addressed by methods like VAMP [29], limiting its applicability to highly noisy data. Additionally, the performance of NN-ResDMD is sensitive to hyperparameter tuning, including network architecture, dictionary size, and training criteria, which can require significant effort to optimize.
Koopman eigenpairs provide unique perspectives into the interpretation of nonlinear dynamical mechanisms, and feedforward neural networks (FNNs) represent an initial step in learning spectral properties directly from data. In recent years, various deep neural network structures have been employed to learn the Koopman representations with different optimization targets other than the spectral residuals (e.g. [30, 21, 29, 31, 22, 32, 33, 34](see Appendix section A.4 for a comparison with the VAMP framework). With our approach, we demonstrate that even basic architectures can achieve significant improvements in Koopman operator estimation by using the spectral residual loss. Therefore, future work could focus on refining neural network architectures to enhance the accuracy and efficiency of Koopman eigenpair estimation. One promising direction is the incorporation of Physics-Informed Neural Networks (PINNs) and Physics-Informed Neural Operators (PINOs), which integrate physical laws directly into the learning process. This integration will ensure that the resulting Koopman eigenfunctions align with known physical constraint, avoid overfitting and faciliatates generalization. Indeed, the integration of PINNs and PINOs with the Koopman framework has the potential to serve as a powerful bridge between data-driven and model-driven approaches, offering enhanced insights into complex systems and enabling more robust temporal evolution predictions.
@@ Acknowledgments
Authors are grateful to Matthew Colbrook for valuable guidance on ResDMD technical part and providing fluid data, to Bin Han for discussion on computational issues, to Michel Beserve and Rory Bufacchi for proofreading and to Turishcheva Polina for providing permission of using the Sensorium 2023 dataset.
@@ Appendix A Appendix
A.1 Calculation steps for 3.3
Here we are going to show how squared relative residual implies (A.1) and then implies (3.3). Consider $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i}$ with $\|\phi\|_{2}=1$, then
@@
[…]
@@
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\|\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\|_{2}^{2}$
@@
$=\frac{d}{dK}\operatorname{tr}\bigg{(}\frac{1}{m}\sum_{i=1}^{N_{K}}\mathbf{v}_{i}^{*}\bigg{[}\Psi_{Y}^{*}\Psi_{Y}-(\Psi_{X}^{*}\Psi_{Y})^{*}K$
$\qquad-K^{*}(\Psi_{X}^{*}\Psi_{Y})+K^{*}(\Psi_{X}^{*}\Psi_{X})K\bigg{]}\mathbf{v}_{i}\bigg{)}$
@@
A.3 Discussion On Convergence
To understand how neural networks enhance NN-ResDMD, it is important to introduce Barron space [35, 36, 23, 37]. Barron space characterizes functions efficiently approximated by two-layer neural networks, which is central to NN-ResDMD. By leveraging networks that approximate functions within this space, NN-ResDMD can flexibly optimize the dictionary functions for Koopman operator approximation, making it highly effective for complex, high-dimensional systems.
@@ A.3 Discussion On Convergence
where $P_{f}$ is the set of distributions for which $f$ can be represented. This framework provides a basis for analyzing approximation errors in neural networks.
@@
$\|f-f_{r}\|_{L^{2}}\lesssim\frac{\|f\|_{\mathcal{B}}}{\sqrt{r}}.$
This result implies that the approximation error decreases at a rate of $1/\sqrt{r}$ as the number of neurons $r$ increases, with the constant $\|f\|_{\mathcal{B}}$ reflecting the complexity of the function $f$ within the Barron space.
Now, consider a Barron space $\mathcal{B}$ which is dense in $L^{2}(\Omega,\mu)$ and a projected Koopman operator $\mathcal{K}_{N_{K}}:\mathcal{B}_{N_{K}}\to L^{2}(\Omega,\mu)$ where $\mathcal{B}_{N_{K}}\subseteq\mathcal{B}$ is a $N_{K}$-dimensional subspace spanned by some dictionary $\mathbf{\Psi}=\{\psi_{i}\}_{i=1}^{N_{K}}$. According to Theorem A.1, we can have a well-trained dictionary that almost spans $\mathcal{B}_{N_{K}}$, i.e., given $\epsilon>0$, we can always obtain a dictionary $\mathbf{\Psi}_{r}=\{\psi_{r,i}\}_{i=1}^{N_{K}}$ such that $\sum_{i=1}^{N_{K}}\|\psi_{r,i}-\psi_{i}\|^{2}_{2}<\epsilon$.
@@
A.4 Highlights of NN-ResDMD compared with typical existing neural network-based Koopman framework

### v3

@@ 4.3 Neural dynamics identification in mice visual cortex
Since ResKoopNet directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics. To evaluate how effectively ResKoopNet reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel-ResDMD. These methods are selected as representative approaches for handling high-dimensional data.
@@
The dataset is part of the open dataset on mice from the competition “Sensorium 2023" [39, 40]. In the experiments, mice viewed natural videos while their neural signals were recorded via calcium imaging in the primary visual cortex, reflecting the activity of thousands of neurons. Here, we focus on the state partitioning of neural signals. Specifically, in each mouse, six video stimuli were repeatedly shown, creating ideal conditions to define brain states. Neural activity during repeated trials with the same stimuli is assumed to reflect the same underlying dynamic system, enabling Koopman decomposition methods to uncover and separate these brain states.
@@
We applied ResKoopNet and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel-ResDMD) to this dataset, using different implementations and Koopman subspace dimensions. For ResKoopNet, we trained dictionaries on all snapshots from each mouse to avoid overfitting. We reduced the data to 24 dimensions via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions. The neural network we use here consists of 3 hidden layers, each with 200 neurons. The hyperparameter scanning results are included in Appendix Figure 17, demonstrating that our parameter choice is within the robust range. The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations. For Hankel-DMD, we built a Hankel matrix with a delay of 50, producing 50 eigenfunctions per trial. In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions. For Kernel-ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on Colbrook et al. [9]. See Appendix A.9.4 for method details and Appendix A.10 for dictionary size justification. These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 14A, and Appendix Figure 15A, are compared to ground truth trial identities.
@@
The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli. To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, which distinguishes them from other states. If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others. This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli. Note that averaged trial differences are even visibly clear for the ResKoopNet case. It is worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 1112). However, this dataset provides an ideal benchmark with clear ground truth labels, enabling accurate comparison of Koopman eigenfunction estimation methods. Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, offering promising directions for future research.
@@
We use Multi-dimensional Scaling (MDS) to visualize how these eigenfunction-based features cluster according to ground truth states. MDS reduces data dimensionality based on similarities, making it ideal for visualizing clustering performance. While Uniform Manifold Approximation and Projection (UMAP) and t-distributed Stochastic Neighbor Embedding (t-SNE) are alternative methods, we show MDS results in 2D space (Figure 6B-E), with similar results for UMAP and t-SNE in Appendix Figure 13, Appendix Figure 14C, D and Appendix Figure 15C, D.
@@
The 2D MDS visualization reveals a clear separation of features for all 5 mice using ResKoopNet (Figure 6B), whereas no other method shows clear clustering (Figure 6C-E, Appendix Figure 14B, Appendix Figure 15B). To quantify this clustering, we calculate the Davies-Bouldin index (DBI), a measure of clustering quality that assesses how compact and well-separated the clusters are. A lower DBI indicates more compact clusters that are farther apart from each other, which corresponds to better clustering. The DBI is significantly lower for ResKoopNet (Figure 6F), suggesting that it captures the latent dynamic structure more effectively than all three other methods. Similar clustering patterns are confirmed with UMAP and t-SNE (Appendix Figure 16).
@@ 5 Conclusion and future work
In this paper, we introduced ResKoopNet, a novel approach that minimizes spectral residuals to accurately approximate Koopman operators for complex dynamical systems. This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high fidelity. Our experiments across physical and biological systems demonstrate that ResKoopNet significantly outperforms existing methods in identifying coherent structures and uncovering latent dynamics, particularly in high-dimensional applications.
Despite these strengths, ResKoopNet has several limitations. First, the use of neural networks increases computational cost compared to classical algorithms (see Appendix A.6). Second, unlike stochastic approaches such as VAMP [25] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison). In addition, its performance can vary depending on the neural architecture and training configuration.
Many physical laws, such as conservation principles or symmetry constraints, are inherently reflected in the spectral properties of dynamical systems. Incorporating such physical structure into the neural network architecture may therefore enhance the learning of Koopman spectral information. Future work could explore physics-informed designs that embed prior knowledge directly into the learning process, allowing the network to better align with the system’s underlying dynamics. This integration may improve both the interpretability and accuracy of the extracted Koopman representations.
@@ Acknowledgments
Authors are grateful to Matthew Colbrook for valuable guidance on ResDMD technical part and providing fluid data, to Bin Han for discussion on computational issues, to Rouslan Krechetnikov for discussion on turbulence analysis, to Michel Beserve and Rory Bufacchi for proofreading and to Turishcheva Polina for providing permission of using the Sensorium 2023 dataset.
@@ Appendix A Appendix
A.1 Source Code
For reproducibility, the source code will be available at the following link.
A.2 Calculation steps for Eq. (3.3)
Here we are going to show how spectral residual implies (3.3).
Consider $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}\mathbf{v}_{i}$ with $\|\phi\|_{\mu}=1$, then
@@ A.1 Source Code
$=\sum_{i,j=1}^{N_{K}}\bar{\mathbf{v}}_{i}\left[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}\right]\mathbf{v}_{j}$
@@
$=\sum_{i=1}^{N_{K}}\frac{1}{m}\|\Psi_{Y}\mathbf{v}_{i}-\Psi_{X}K\mathbf{v}_{i}\|_{\mu}^{2}$
@@
$=\frac{d}{dK}\operatorname{tr}\bigg{(}\frac{1}{m}\sum_{i=1}^{N_{K}}\mathbf{v}_{i}^{*}\bigg{[}\Psi_{Y}^{*}\Psi_{Y}-(\Psi_{X}^{*}\Psi_{Y})^{*}K-K^{*}(\Psi_{X}^{*}\Psi_{Y})+K^{*}(\Psi_{X}^{*}\Psi_{X})K\bigg{]}\mathbf{v}_{i}\bigg{)}$
@@
A.4 Discussion on convergence analysis
To understand how neural networks enhance ResKoopNet and build up the theoretical framework of convergence, it is important to first introduce Barron space [31, 10, 14, 2]. Barron space characterizes functions efficiently approximated by two-layer neural networks, which is central to ResKoopNet. By leveraging networks that approximate functions within this space, ResKoopNet can flexibly optimize the dictionary functions for Koopman operator approximation, making it highly effective for complex, high-dimensional systems.
@@ A.4 Discussion on convergence analysis
where $P_{f}$ is the set of distributions for which $f$ can be represented. This framework provides a foundation for analyzing approximation errors in neural networks.
@@
$\|f-f_{r}\|_{\mu}\lesssim\frac{\|f\|_{\mathcal{B}}}{\sqrt{r}}.$
This implies an approximation error decreasing at a rate of $O(1/\sqrt{r})$, where $r$ is the number of neurons. In ResKoopNet, the dictionary $\Psi(x;\theta)=\{\psi_{i}(x;\theta)\}_{i=1}^{N_{K}}$ is parameterized by a neural network with parameters $\theta$. Assuming the true dictionary functions $\psi_{i}\in\mathcal{B}$, Theorem A.2 ensures that $\Psi(x;\theta)$ can approximate the optimal dictionary spanning the Koopman invariant subspace $\mathcal{B}_{N_{K}}\subset\mathcal{F}$ with error $\epsilon>0$, provided $r$ is sufficiently large.
Remark A.3.
Notice that, the “two-layer neural network" in the Theorem A.2 statement refers to a hidder layer + an output layer, which is the most standard and general setting. Our implementation uses three hidden layers. This does not invalidate the result, as deeper networks can achieve at least the same approximation power [31].
We want to show two convergence results here: (1) $\Psi(x;\theta)$ approaches the true invariant subspace of $\mathcal{K}$; (2) The eigenpairs $(\lambda_{i},\phi_{i})$ and pseudospectrum approximate $\mathcal{K}$’s true spectrum as $J(\theta)\to 0$.
Assumption A.4.
To formalize convergence, we make the following assumptions:
  1. (a)
    The optimal dictionary functions $\{\psi_{i}^{*}\}_{i=1}^{N_{K}}$ spanning $\mathcal{K}$’s invariant subspace lie in $\mathcal{B}$.
  2. (b)
    The loss $J(\theta)$ is Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
Now, consider a Barron space $\mathcal{B}$ which is dense in $\mathcal{F}$. Given $N_{K}$ fixed, let $\mathcal{B}_{N_{K}}=\text{span}\{\psi_{i}^{*}\}_{i=1}^{N_{K}}$ be the true invariant subspace. By Theorem A.2, for each $\psi_{i}^{*}$, there exists a neural network approximation $\psi_{i}(x;\theta_{r})$ with $r$ neurons such that:
$\|\psi_{i}^{*}-\psi_{i}(\cdot;\theta_{r})\|_{\mu}\leq\frac{\|\psi_{i}^{*}\|_{\mathcal{B}}}{\sqrt{r}}.$
The total dictionary error is:
$\|\Psi^{*}-\Psi(\cdot;\theta_{r})\|_{F}^{2}=\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}-\psi_{i}(\cdot;\theta_{r})\|_{\mu}^{2}\leq\frac{1}{r}\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}\|_{\mathcal{B}}^{2}=\frac{C_{N_{K}}}{r},$
where $C_{N_{K}}=\sum_{i=1}^{N_{K}}\|\psi_{i}^{*}\|_{\mathcal{B}}^{2}$ is finite under Assumption A.4(a). As $r\to\infty$, $\Psi(x;\theta_{r})\to\Psi^{*}$ in the Frobenius norm, which ensures the dictionary approximated by neural network can represent $\mathcal{B}_{N_{K}}$.
Algorithm 1 updates $\theta$ via stochastic gradient descent (SGD) with step size $\delta$ and computes $\tilde{K}(\theta_{n})$ iteratively until $J(\theta_{n})<\epsilon$ where $\theta_{n}$ represents $n$-th iteration of parameters $\theta$. For a Lipschitz continuous $J(\theta)$ with constant $L$ (by Assumption A.4(b)) and a strongly convex region near the optimum $\theta^{*}$ (assumed locally for simplicity), SGD converges at a rate of $O(1/n)$ in expectation [4]:
$\mathbb{E}[J(\theta_{n})-J(\theta^{*})]\leq\frac{L}{2\eta n},$
where $\eta$ is the strong convexity constant and $n$ is the iteration number. In practice, $J(\theta)$ is non-convex due to the neural network, and the alternating update with $\tilde{K}(\theta)$ stabilizes the process. As iteration step $n\to\infty$ and amount of data points $m\to\infty$, $J(\theta_{n})\to 0$, which implies $\widehat{\text{res}}(\lambda_{i},\phi_{i})\to 0$ for all $i$.
With $J(\theta)\to 0$, we now assess the spectral error: if $\widehat{\text{res}}(\lambda_{i},\phi_{i})<\epsilon$, then $\|\mathcal{K}\phi_{i}-\lambda_{i}\phi_{i}\|_{\mu}/\|\phi_{i}\|_{\mu}<\sqrt{\epsilon}$, indicating $\lambda_{i}$ and $\phi_{i}$ are approximate eigenpairs of $\mathcal{K}$ converges to $\mathcal{K}$’s true spectrum as $N_{K}\to\infty$ and $\epsilon\to 0$, as pointed out in [8, Theorem B.1]. Uniform convergence on compact subsets of $\mathbb{C}$ follows from the density of $\mathcal{B}_{N_{K}}$ in $\mathcal{F}$ and Dini’s theorem, as pointed out in Colbrook and Townsend [8, Lemma B.1].
The non-convexity and system complexity may slow the spectral approximation in practice. High-order convergence (e.g., polynomial) could arise for smooth dynamics (See Colbrook and Townsend [8, Theorem 3.1] for more details), which is useful for further study.
Remark A.5.
The computational cost (Appendix A.6) scales with $r$, $t$, and $m$, which trades off with accuracy. Adaptive selection of $r$ and early stopping when $J(\theta)<\epsilon$ can optimize this balance.
@@
A.5 Highlights of ResKoopNet compared with typical existing neural network-based Koopman framework